### v1

# coverforest: Conformal Predictions with Random Forest in Python

Abstract
Conformal prediction provides a framework for uncertainty quantification, specifically in the forms of prediction intervals and sets with distribution-free guaranteed coverage. While recent cross-conformal techniques such as CV+ and Jackknife+-after-bootstrap achieve better data efficiency than traditional split conformal methods, they incur substantial computational costs due to required pairwise comparisons between training and test samples’ out-of-bag scores. Observing that these methods naturally extend from ensemble models, particularly random forests, we leverage existing optimized random forest implementations to enable efficient cross-conformal predictions.
We present coverforest, a Python package that implements efficient conformal prediction methods specifically optimized for random forests. coverforest  supports both regression and classification tasks through various conformal prediction methods, including split conformal, CV+, Jackknife+-after-bootstrap, and adaptive prediction sets. Our package leverages parallel computing and Cython optimizations to speed up out-of-bag calculations. Our experiments demonstrate that coverforest’s predictions achieve the desired level of coverage. In addition, its training and prediction times can be faster than an existing implementation by 2–9 times. The source code for the coverforest is hosted on GitHub at https://github.com/donlapark/coverforest.
Keywords: conformal prediction, cross-validation, jackknife, bootstrap, prediction sets
1 Introduction
Classification is a fundamental machine learning task with applications ranging from medical diagnosis to customer segmentation and computer vision. While numerous classification algorithms exist in the literature, most implementations focus solely on point predictions. However, in many real-world applications, particularly those involving critical decision-making, understanding the uncertainty associated with these predictions becomes crucial. For this reason, conformal prediction techniques have recently garnered significant interest, as they leverage data exchangeability to produce prediction sets/intervals that contain the true output with a guaranteed probability above a specified level.
Recent conformal prediction methods like CV+ and Jackknife+-after-bootstrap (J+ab) rely on out-of-fold and out-of-bag calculations—techniques extensively studied in ensemble learning, particularly in random forests [1]. Years of development in open-source machine learning packages have led to highly optimized out-of-bag calculations in random forests. The scikit-learn’s implementation, for instance, leverages parallel computing to speed up the calculations. Even though these implementations are only for in-sample out-of-bag calculations, we can adapt the technique for train-test cross-conformity score calculations, enabling faster conformal predictions with CV+ and J+ab.
Building on this idea, we introduce coverforest, a Python package for conformal predictions using random forest. coverforest implements several recent advances in conformal prediction, including split conformal, CV+ and J+ab for regression, and adaptive prediction set (APS) with split conformal, CV+, and J+ab for classification. Random forest is our model of choice for three reasons 1) they have consistently demonstrated good performance on tabular data [2], 2) they can be trained with missing data, and 3) trees are very fast to train in parallel, making them attractive for cross-conformal predictions like CV+ and J+ab.
In our implementation, trees are fitted and queried for predictions in parallel using the joblib [3] package in Python, and the cross-conformity scores are calculated in Cython [4]. Experimental results show that coverforest’s conformal training and prediction of random forests are faster than an existing implementation by 2–9 times faster.
2 Background on Conformal Predictions
Let $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ be training data for a predictive model $\hat{f}$ and $(X_{n+1},Y_{n+1})$ is a new test point, where $X_{i}$’s lie in an input space $\mathcal{X}$ and $Y_{i}$’s lie in an output space $\mathcal{Y}$. Without loss of generality, we assume that $\mathcal{Y}=\mathbb{R}$ for a regression task and $\mathcal{Y}=\{1,\ldots,C\}$ for a classification task. A prediction set for $Y_{n+1}$ is a set of outputs $\mathcal{C}=\mathcal{C}(X_{i})\subseteq\mathcal{Y}$ that might be input-dependent. For regression tasks prediction sets are usually restricted to intervals, in which case they are commonly referred to as prediction intervals.
In a conformal prediction task, we specify a miscoverage rate $\alpha\in(0,1)$, with which we aim to construct a prediction set $\mathcal{C}$ that has probability at least $1-\alpha$ to contain the true class:
$P(Y_{n+1}\in\mathcal{C}(X_{n+1}))\geq 1-\alpha.$ (1)
Classical results in conformal prediction literature typically assume that the random variables $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ and $(X_{n+1},Y_{n+1})$ are either i.i.d. or exchangable, that is, their joint distribution are invariant under permutations. Below, we summarize relevant conformal prediction methods that provide prediction sets with coverage guarantees under one of these assumptions.
Regression
  1. 1.
    Split Conformal [5]. First, we split the data into training and calibration sets, we fit a predictive model $\hat{f}$ on the training data. The method then computes nonconformity scores $R_{i}=|Y_{i}-\hat{f}(X_{i})|$ for each point in the calibration set. For the new test point $X_{n+1}$, the prediction interval is constructed as:
    $\mathcal{C}(X_{n+1})=[\hat{f}(X_{n+1})-q_{\alpha},\hat{f}(X_{n+1})+q_{\alpha}],$
    where $q_{\alpha}$ is the $\lceil(1-\alpha)(n+1)\rceil$-th smallest value of the calibration scores.
  2. 2.
    Cross-Validation+ (CV+) [6]. Split the data into $K$ folds. For each fold $\mathcal{I}_{k}$, we fit a model $\hat{f}^{(-k)}$ on all data except fold $k$ and compute residuals $R_{i}=|Y_{i}-\hat{f}^{(-k)}(X_{i})|$ for all $i\in\mathcal{I}_{k}$. The final prediction interval for a new point is constructed as:
    $\mathcal{C}(X_{n+1})=[q^{-}_{\alpha}\{\hat{f}^{-S(i)}(X_{n+1})-R_{i}\},q^{+}_{\alpha}\{\hat{f}^{-S(i)}(X_{n+1})+R_{i}\}],$
    where $S(i)=k$ if $i\in\mathcal{I}_{k}$, $q^{+}_{\alpha}\{a_{i}\}$ is the $\lceil(1-\alpha)(n+1)\rceil$-th smallest value of $a_{i}$’s and $q^{-}_{\alpha}\{a_{i}\}$ is the $\lfloor\alpha(n+1)\rfloor$-th smallest value of $a_{i}$’s.
  3. 3.
    Jackknife+-after-Bootstrap (J+ab) [7]. This method is similar to CV+ but instead of splitting data into $K$-folds, we sample with or without replacement from the original training set multiple times, and fit a model on each of the subsamples. During prediction, we let $\hat{f}^{(-i)}(X_{i})$ be the aggregated predictions from all models that were fitted on the subsamples that do not contain $(X_{i},Y_{i})$, from which we compute the residuals $R_{i}=|Y_{i}-\hat{f}^{(-i)}(X_{i})|$. The final prediction interval for a new point is then
    $\mathcal{C}(X_{n+1})=[q^{-}_{\alpha}\{\hat{f}^{(-i)}(X_{n+1})-R_{i}\},q^{+}_{\alpha}\{\hat{f}^{(-i)}(X_{n+1})+R_{i}\}],$ (2)
    where $q^{-}_{\alpha}$ and $q^{+}_{\alpha}$ are defined as in CV+. However, this interval with a fixed number of bootstrap samplings $B$ does not achieve a finite-sample coverage guarantee, due to the asymmetry between training and test samples; to see this, observe that each test sample is excluded from exactly $B$ bootstraps, while each training sample is typically excluded from $<B$ bootstraps. This issue is resolved by making $B$ a random variable—by resampling $B$ from a Binomial distribution:
    $\operatorname{Binomial}(\tilde{B},(1-\frac{1}{n+1})^{m}),$ (3)
    where $\tilde{B}$ is a fixed initial value, $n$ is the number of training samples and $m$ is the number of bootstrapped samples. One can show that the interval in (2) with the randomized $B$ restores the symmetry between training and test samples [7].
Classification
  1. 1.
    Adaptive Prediction Sets (APS) [8]. This method proposes the cumulative probability as the conformity score: given the predicted probabilities of a point $(x,y)$ in descending order $\hat{\pi}_{1}(x)\geq\hat{\pi}_{2}(x)\geq\ldots\geq\hat{\pi}_{C}(x)$, the score function is the (randomized) generalized inverse quantile:
    $E(x,y)=\hat{\pi}_{1}(x)+\ldots+\hat{\pi}_{y-1}(x)+u\cdot\hat{\pi}_{y}(x),\quad u\sim\operatorname{Uniform}(0,1).$ (4)
    In other words, the score is chosen randomly between the cumulative probabilities of $y$ and $y-1$. Without this randomization, the prediction sets would generally cover more than the desired level of ($1-\alpha$)if $E(x,y)$ is defined up to $\hat{\pi}_{y}$, and less if it is defined up to $\hat{\pi}_{y-1}$.
    For the split conformal, we let $q_{\alpha}$ be the $\lceil(1-\alpha)(n+1)\rceil$-th smallest value of the calibration scores. The prediction set for a new point $X_{n+1}$ is then:
    $\mathcal{C}(X_{n+1})=\left\{1,\ldots,k:k=\inf_{y\in\{1,\ldots,C\}}E(X_{n+1},y)\geq q_{\alpha}\right\}.$
    There is also a CV+ variant of this method: First, we split the data into $K$ folds. For each fold $\mathcal{I}_{k}$, we fit a model $\hat{\pi}^{(-k)}$ on all data except fold $k$ and then calculate the score $E(X_{i},Y_{i},\hat{\pi}^{(-S(i))})$ where $S(i)=k$ if $i\in\mathcal{I}_{k}$. The prediction set for a new point $X_{n+1}$ is then all $y$ such that less that $1-\alpha$ fraction of the training set has their scores less than the score of $y$ under the same model, in other words, $E(X_{i},Y_{i},\hat{\pi}^{(-S(i))})<E(X_{n+1},y,\hat{\pi}^{(-S(i))})$.
  2. 2.
    Regularized Adaptive Prediction Sets (RAPS). [9]. This method was proposed for classification with a lot of classes (say 1000+ classes) where the ordering of the predicted probabilities does not reflect the true order. To this end, the authors propose to add to the APS’s score function a regularization term that penalizes classes with small predicted probabilities (assuming $\hat{\pi}_{1}(x)\geq\hat{\pi}_{2}(x)\geq\ldots\geq\hat{\pi}_{C}(x)$ as before):
    $E_{k,\lambda}(x,y)=E(x,y)+\lambda\cdot\max(0,y-k).$ (5)
    As with APS, we can use this score function for split conformal or CV+.
  3. 3.
    Jackknife+-after-Bootstrap (J+ab). This method extends Jackknife+after-bootstrap from regression to classification by using APS or RAPS score functions. Since the proof of the coverage lower bound applies to any exchangeable conformity scores, the lower bound remains valid when using these classification scores.
Method Theoretical Practical Training Prediction
Coverage Coverage Cost Cost
Regression
Split Conformal $1-\alpha$ $\approx 1-\alpha$ $O(Tdn\log n)$ $O(Tn_{\operatorname{test}}\log n)$
CV+ $1-2\alpha-\epsilon_{K,n}$ $\gtrsim 1-\alpha$ $O(KTdn\log n)$ $O(KTn_{\operatorname{test}}\log n)$
J+ab $1-2\alpha$ $\gtrsim 1-\alpha$ $O(Tdn\log n)$ $O(n*n_{\operatorname{test}})$
Classification
APS + Split $1-\alpha$ $\approx 1-\alpha$ $O(Tdn\log n)$ $O(Tn_{\operatorname{test}}\log n)$
APS + CV+ $1-2\alpha-\epsilon_{K,n}$ $\gtrsim 1-\alpha$ $O(KTdn\log n)$ $O(Cn*n_{\operatorname{test}})$
APS + J+ab $1-2\alpha$ $\gtrsim 1-\alpha$ $O(Tdn\log n)$ $O(Cn*n_{\operatorname{test}})$
Table 1: Comparison of coverage guarantees and computational costs for different conformal random forest methods. Here, the notation $\gtrsim 1-\alpha$ means “sometimes worse than $1-\alpha$”, $n$ is the number of training samples, $n_{\operatorname{test}}$ is the number of test samples, $T$ is the number of trees, $C$ is the number of classes, $K$ is the number of folds in CV+, and $\epsilon_{K,n}=\min\left\{2(1-1/K)/(n/K+1),(1-K/n)/(K+1)\right\}$. We assume that $T\ll n$.
In Table 1, we compare coverage guarantees and computational costs across different conformal prediction methods for random forests. While CV+ and J+ab theoretically guarantee coverages of $\approx 1-2\alpha$, they typically achieve $\geq 1-\alpha$ coverage in practice. The methods with the APS score, however, have notably long prediction times due to the pairwise comparisons between train and test conformity scores across all classes.
2.1 Related work
The coverforest package implement conformal predictions [10, 5, 11, 12, 13] using random forest [1]. For similar work, conformity scores of bootstrapped regression based on iterative reweighted least squares is briefly discussed in [5]. For random forest classification, three conformity scores were proposed in [14], though their coverage guarantee held only with infinite data.
Nonconformity scores based on out-of-bag (OOB) predictions were introduced in [15]. The experiments revealed that these scores yield smaller prediction intervals compared to neural networks and $k$-nearest neighbors. Later, it is confirmed in [16] that these intervals provide asymptotic coverage guarantee, which is only attained with infinitely many data. Further works explored various aspects of out-of-bag calibration for conformal predictions [17, 18, 19]. coverforest’s implementation of Jackknife+-after-bootstrap uses a randomized number of bootstraps in order to achieve a finite-sample coverage guarantee, which is proved in [7].
A conformal regression method related to CV+ is Jackknife+ [6] where the regression model is fitted on leave-one-out datasets. It is a special case of CV+ where each fold only consists of a single training point, and the number of folds equals the number of training samples. coverforest can be used with the Jackknife+ method by specifying method="cv" and cv equals the number of training samples. Another method is the conformalized quantile regression (CQR) which constructs prediction intervals based on classical quantile regression. Implementing CQR is one of our future plans where we can incorporate existing implementation of conformal quantile random forest [20] such as quantile-forest [21].
For classification, given a model that outputs probability estimates $\hat{\pi}_{j}(x)\approx P(Y=j|X=x)$, a basic conformal prediction method can be devised based on a score function $s(x,y)=1-\hat{\pi}_{y}(x)$. The prediction set for a new point $X_{n+1}$ is then given by $\mathcal{C}(X_{n+1})=\left\{y:s(X_{n+1},y)\leq q_{\alpha}\right\}$, where $q_{\alpha}$ is the $\lceil(1-\alpha)(n+1)\rceil$-th smallest value of the calibration scores [22, 5]. However, this method tend to outputs an empty set when the model is uncertain about the prediction, for example, $q_{\alpha}=0.2$ and the predicted probabilities of two classes are 0.5 each. This issue can be remedied by instead using the cumulative probability as in the APS method [8]. Our future plans also include implementing multilabel classification with finite-sample coverage guarantee via conformal risk control [23, 24, 25].
3 Implementation
3.1 API description
Figure 1: Class structure and the metrics in coverforest.
coverforest’s provides two main model classes with familiar scikit-learn-style API, with the standard fit() and predict() methods, as shown in Figure 1. The most important parameters of both classes are:
  • n_estimators: The number of sub-estimators in the forest
  • method: The conformal prediction methods, which can be one of cv (CV+), bootstrap (Jackknife+-after-bootstrap), and split, (split conformal)
  • cv: The number of folds in the CV+ method
There are also several parameters that are only specific to APS and RAPS:
  • k_init_ and lambda_init_: The regularization parameters in RAPS. The APS score is equivalent to setting k_init_=0 and lambda_init_=0. Sutomatic searching using the method described in [9] is also available by setting k_init_="auto" and lambda_init_="auto"
  • allow_empty_set: whether to allow empty set predictions
  • randomized: whether to apply the APS’s randomization scheme to include an additional class in prediction sets as described in Section 2.
There are only one parameter that is specific to the Jackknife+-after-bootstrap method:
  • resample_n_estimators: whether to draw the number of trees from the Binomial distribution (3), in which case n_estimators is taken as the first parameter $\tilde{B}$ of the distribution.
We now go over the methods in both model classes.
The fit() method fits each tree on the training data. The trees are fitted in parallel via a helper class joblib.Parallel() [3]. The fitted trees are then used to calculate for each training sample the out-of-box (OOB) predictions, which are then used to calculate the calibration scores. During fitting, the method maintains $\mathcal{T}_{i}$ for each sample $(X_{i},Y_{i})$, where $\mathcal{T}_{i}$ is the collection of trees not fitted on this particular sample.
The predict() method first queries predictions from all trees for any given test sample $(X_{n+1},Y_{n+1})$, where the queries are performed in parallel via joblib.Parallel(). Subsequently, for each training instance $i$, it aggregates predictions exclusively from trees in $\mathcal{T}_{i}$, converting these into a conformity score. This score is then compared with the calibration scores of $(X_{i},Y_{i})$ to produce the final predictions.
For moderately large training and test sets, calculating the conformity scores across all pairs of training and test samples can be quite expensive. To improve performance, the calculations have been optimized in Cython with parallel processing enabled. Users can specify the number of threads through the num_threads parameter in the predict() method.
The CoverForestClassifier includes the search_k_and_lambda() method for automatic parameter optimization. This method determines optimal values for $k$ and $\lambda$ according to the approach detailed in [9].
coverforest also has a metric module that provides convenient functions for evaluating the prediction sets/intervals by measuring their average size and empirical coverage (e.g. the proportion of prediction sets/intervals that contains the true class/value).
3.2 Source code
The source code for the coverforest is hosted on GitHub at https://github.com/donlapark/coverforest. The code follows the PEP8 guidelines, and all docstrings of publicly exposed classes comply with the NumPy docstring style. Both CoverForestClassifier and CoverForestRegressor pass the comprehensive unit tests provided by scikit-learn. The documentation includes installation guides, user manual, and API reference, generated using sphinx and available online at https://donlapark.github.io/coverforest. The package is BSD-licensed and requires only scikit-learn version 1.6.1 or greater.
4 Examples
We demonstrate how to use coverforest on a toy dataset for set predictions.
Conformal predictions with CV+ and RAPS score
We train a conformal random forest using CV+ with the RAPS score. To use the CV+ method, initialize a CoverForestClassifier model with method="cv". Here, we set cv=5 to split the data into five folds. As a result, five random forests will be fitted on the five combinations of 4-folds. We also specify that each forest has 200 trees by setting n_estimators=200. {minted}[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=]python from coverforest import CoverForestClassifier from coverforest.metrics import average_set_size_loss from coverforest.metrics import classification_coverage_score from sklearn.datasets import load_digits from sklearn.model_selection import train_test_split
X, y = load_digits(return_X_y=True) X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2)
clf = CoverForestClassifier(n_estimators=200, method="cv", cv=5)
Training and making predictions
We train the random forest by calling the fit() method, and make a prediction by calling the predict() method. During prediction, we choose the miscoverage rate at $0.05$ by specifying alpha=0.05. The output of the predict() method is a tuple: The first item contains the single class predictions by the random forest, and the second item contains the set predictions. {minted}[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=]python clf.fit(X_train, y_train) y_pred, y_sets = clf.predict(X_test, alpha=0.05) # y_pred[:5] = [2, 8, 2, 6, 6] # y_sets[:5] = [array([2]), array([2, 5, 8]), array([2]), # array([5, 6]), array([6])]
Computing average set size and coverage
We may use the provided metrics functions to evaluate the set predictions against the true classes. {minted}[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=]python avg_set_size = average_set_size_loss(y_test, y_sets) coverage = classification_coverage_score(y_test, y_sets) # avg_set_size = 1.589 # coverage = 0.964 The small average set size indicates that the model can make predictions with high confidence. The coverage tells us that $96.4\%$ of the prediction sets contain the true classes, which is above our specified level of $95\%$.
Conformal predictions using Jackknife+-after-bootstrap with the APS score
In the second example, we train a conformal random forest using Jackknife+-after-bootstrap (J+ab) with the APS score. To use J+ab, initialize a CoverForestClassifier model with method="bootstrap". To use the APS method, set k_init=0 and lambda_init=0. In contrast to CV+ which has multiple forests, J+ab has only a single forest. Here, we specify that this forest has 400 trees. {minted}[ frame=lines, framesep=2mm, baselinestretch=1.2, fontsize=]python clf = CoverForestClassifier( n_estimators=400, method="bootstrap", k_init=0, lambda_init=0 ) clf.fit(X_train, y_train) y_pred, y_sets = clf.predict(X_test, alpha=0.05) avg_set_size = average_set_size_loss(y_test, y_sets) coverage = classification_coverage_score(y_test, y_sets) # y_pred[:5] = [2, 8, 2, 6, 6] # y_sets[:5] = [array([2]), array([2, 5, 8, 9]), # array([2]), array([6]), array([6])] # avg_set_size = 1.761 # coverage = 0.968 Compared to the CV+ method, J+ab produces prediction sets with a larger average with practically the same coverage.
5 Experiments and Discussions
To validate our implementation, we benchmark coverforest’s three methods: CV+, Jackknife+-after-bootstrap (J+ab), and split conformal on conformal classfication and regression tasks.
5.1 Classification experiment
We evaluate the predictions and run times of CoverForestClassifier on four classification datasets. See Table 2 for the datasets’ characteristics.
To evaluate a model on each dataset, we split the data with 80% as the training set and 20% as the test set; then we fit the models on the training set with the target miscoverage level $\alpha\in\{0.05,0.1,0.2\}$. We repeat the experiment 50 times for each method and compare the empirical coverage (the proportion of prediction sets that contain the true class), and average set size. For fair comparisons, we balance the number of trees between the three methods so that they have the same predictive power. This involves ensuring that each training sample is be used to train the same number of trees across all these models. For CV+, the data is split into 10 folds, and we fit 10 forests with 100 trees each, totaling 1000 trees. Since each tree in CV+ is trained on 90% of the data, we accordingly set both J+ab and split conformal methods to use 900 trees to maintain consistency in the learning capacity.
Instances Classes Features Training set Test set
Mice [26] 1080 8 80 864 216
WineQuality [27] 2000 7 11 1600 400
Myocardial [28] 1700 8 111 1360 240
MNIST [29] 1250 10 784 1000 250
Table 2: The datasets used in the conformal classification experiment.
Figure 2: Coverage probabilities of CoverForestClassifier’s set predictions on the test set, repeated 50 times for each of the CV+, Jackknife+-after-bootstrap (J+ab) and split conformal methods with the APS and RAPS scores.
Empirical coverage
Figure 2 shows the empirical coverages for all methods using both APS (4) and RAPS (5) scores. In all datasets, the median coverages generally align with the nominal levels. What is most surprising is the MNIST datasets, where all methods achieve the coverages way above the nominal levels.
In the Mice and Myocardial datasets, there is no difference between the APS and RAPS scores. However, in the WineQuality and MNIST datasets, methods with the APS score typically achieve higher coverage than those with the RAPS score, with CV+ methods consistently outperforming J+ab in terms of coverage.
Average set size
Figure 3 shows the average set sizes across the methods. In all datasets, split conformal exhibits notably higher variance in set sizes compared to CV+ and J+ab methods. The RAPS score’s regularization effectively reduced set sizes in the Mice, Myocardial, and MNIST datasets, with the most pronounced reduction observed in MNIST. This substantial reduction in MNIST can be attributed to the visual similarity between certain digit pairs (e.g., 4/9, 3/8, 1/7), which causes the model to distribute non-negligible probabilities across multiple classes even for clear instances. In both WineQuality and MNIST datasets, J+ab consistently produced smaller prediction sets than CV+, corresponding to its lower coverage rates observed in these datasets.
Figure 3: Average size of CoverForestClassifier’s set predictions on the test set, repeated 50 times for each of the CV+, Jackknife+-after-bootstrap (J+ab) and split conformal methods with the APS and RAPS scores.
Run time comparison with MAPIE
In this experiment, we compare the run time of CoverForestClassifier against MAPIE [30], a Python package that implement wrapper classes that enable conformal prediction on any scikit-learn models. It is currently the only package that provides the CV+ and J+ab methods with the APS and RAPS scores (though J+ab with APS and RAPS for conformal classification has not been implemented). MAPIE’s development emphasizes versatility across all machine learning models, whereas coverforest’s development focuses on efficient random forest implementations.
For coverforest and MAPIE, we train a random forest and make conformal predictions using CV+, J+ab and split conformal methods, all with the APS score. We fit the models across the four datasets introduced in Experiment 1. We now use the full WineQuality dataset of 3918 training and 980 test samples, and the MNIST dataset is expanded to 5000 training and 1250 test samples. We run the three methods with the same number of trees as in Experiment 1 (1000 trees for CV+ and 900 trees for J+ab and split conformal). And we fix $\alpha=0.05$. We run each method for 30 repetitions each and report the average training and prediction times.
Figure 4: Training and prediction times of coverforest’s and MAPIE’s implementations of the CV+, Jackknife+-after-bootstrap (J+ab) and split conformal methods with the APS score and $\alpha=0.05$.
Figure 4 compares the run times of both packages (note that at the time of this study, MAPIE had not implemented J+ab with the APS score). It shows that coverforest consistently runs faster than MAPIE in both training and prediction, especially with large datasets like WineQuality and MNIST. For the CV+ method specifically, coverforest’s optimized implementation reduces cross-conformal prediction times to only a few seconds.
5.2 Regression experiment
Instances Output Range Features Training set Test set
Housing [31] 5000 [0.15, 5.0] 8 4000 1000
Concrete [32] 1030 [2.33, 82.6] 8 824 206
Bike [33] 5000 [1, 638] 12 4000 1000
Crime [34] 1994 [0.0, 1.0] 122 1595 399
Table 3: The datasets used in our experiment.
We now evaluate CoverForestRegressor for interval predictions on four regression datasets. The characteristics of the benchmark datasets are detailed in Table 3.
We use the same numbers of trees as in Experiment 1 (1000 trees for CV+ and 900 trees for J+ab and split conformal). For each dataset, we run each method for 50 repetitions and calculate the coverage and average interval length. For benchmarking run times, we use the full Housing dataset (16,512 training, 4,128 test samples) and Bike dataset (13,903 training, 3,476 test samples). We fix $\alpha=0.05$. For each method, we perform training and prediction for 30 repetitions and compare the average run times to those of MAPIE.
Empirical coverage and average interval length
Figure 5: Coverage and average length of CoverForestRegressor’s interval predictions on the test set, repeated 50 times for each of the CV+, Jackknife+-after-bootstrap (J+ab) and split conformal methods with the residuals as the conformity scores.
Figure 5 shows the coverages and average interval lengths of the three methods. Across all datasets, both CV+ and J+ab produce prediction intervals with small variances compared to the split conformal, with J+ab consistently producing intervals of smaller average lengths and smaller coverages than CV+.
Run time comparison with MAPIE
Figure 6 shows the run times of coverforestand MAPIE on conformal regression. While prediction times are similar for both packages, coverforest demonstrates faster training times than MAPIE for all methods and datasets. The run time differences are small on smaller datasets like Concrete due to computational overheads. Conversely, on larger datasets, coverforest’s CV+ and split conformal train substantially faster than MAPIE’s implementations.
Figure 6: Training and prediction times of coverforest’s and MAPIE’s implementations of the CV+, Jackknife+-after-bootstrap (J+ab) and split conformal methods with the residuals as the conformity scores and $\alpha=0.05$.
6 Conclusion
In this work, we introduce coverforest, a lightweight package that bridges the gap between scikit-learn’s optimized random forest implementation and modern conformal prediction techniques. By leveraging both cross-validation+ and Jackknife+-after-bootstrap approaches, coverforest  enables uncertainty quantification through distribution-free prediction intervals and prediction sets.
Our empirical evaluation demonstrates that coverforest  not only achieves the desired coverage guarantees but is also faster than existing implementation of conformal predictions with random forests. Given random forest’s established effectiveness on tabular data, coverforest  provides a practical solution for deployment of conformal prediction in real-world applications. Future work could explore extensions to other ensemble methods and conformal prediction techniques while maintaining the package’s emphasis on simplicity and efficiency.