### v2

# coverforest: Conformal Predictions with Random Forest in Python

Abstract
Conformal prediction provides a framework for uncertainty quantification, specifically in the forms of prediction intervals and sets with distribution-free guaranteed coverage. While recent cross-conformal techniques such as CV+ and Jackknife+-after-bootstrap achieve better data efficiency than traditional split conformal methods, they incur substantial computational costs due to required pairwise comparisons between training and test samples’ out-of-bag scores. Observing that these methods naturally extend from ensemble models, particularly random forests, we leverage existing optimized random forest implementations to enable efficient cross-conformal predictions.
We present coverforest, a Python package that implements efficient conformal prediction methods specifically optimized for random forests. Our package leverages parallel computing and Cython optimizations to accelerate out-of-bag calculations, making it 2–10 times faster than existing implementations while maintaining theoretical coverage guarantees. coverforest supports both regression and classification tasks through various conformal prediction methods, including split conformal, CV+, and Jackknife+-after-bootstrap, as well as adaptive prediction sets for classification. Built upon scikit-learn’s random forest implementation, coverforest provides a practical solution for reliable uncertainty quantification. We demonstrate the package’s performance through empirical evaluations on several benchmark datasets. The source code for the coverforest is hosted on GitHub at https://github.com/donlap/coverforest.
Keywords: conformal prediction, cross-validation, jackknife, bootstrap, prediction sets
1 Introduction
Classification and regression are fundamental machine learning tasks with applications ranging from medical diagnosis to customer segmentation and computer vision. While numerous classification algorithms exist in the literature, most implementations focus solely on point predictions. However, in many real-world applications, particularly those involving critical decision-making, understanding the uncertainty associated with these predictions becomes crucial. For this reason, conformal prediction techniques [1, 2, 3, 4] have recently garnered significant interest, as they leverage data exchangeability to produce prediction sets/intervals that contain the true output with a guaranteed probability above a specified level.
Recent conformal prediction methods like CV+ [5] and J+ab [6] rely on out-of-fold and out-of-bag calculations, techniques frequently used in ensemble learning and particularly in random forests [7]. Years of development in open-source machine learning packages have led to highly optimized out-of-bag calculations in random forests. The scikit-learn’s implementation, for instance, leverages parallel computing to speed up the calculations. Even though these implementations are only for in-sample out-of-bag calculations, we can adapt the technique for train-test cross-conformity score calculations, enabling faster conformal predictions with CV+ and J+ab.
Building on this idea, we introduce coverforest, a Python package for conformal predictions using random forest. coverforest implements several recent advances in conformal prediction, including split conformal [4], CV+ [8] and J+ab [6] for regression, and adaptive prediction set (APS) [5] with split conformal, CV+, and J+ab for classification. Random forest is our model of choice for three reasons 1) they have consistently demonstrated good performance on tabular data [9], 2) they can be trained with missing data [7, 10], and 3) trees are very fast to train in parallel [11], making them attractive for cross-conformal predictions like CV+ and J+ab.
In our implementation, trees are fitted and queried for predictions in parallel using the joblib [12] package in Python, and the cross-conformity scores are calculated in Cython [13].
The following is a summary of coverforest’s contribution:
  1. 1.
    coverforest directly extends scikit-learn’s random forest, which means it also utilizes already well-optimized C+ code and parallel computation. Experimental results show that coverforest’s conformal training and prediction of random forests are 2–10 times faster than an existing implementation.
  2. 2.
    coverforest is designed to follow scikit-learn’s API conventions with the simple fit() and predict() methods. It can also be easily combined with scikit-learn’s preprocessing methods, allowing it to handle various data formats.
  3. 3.
    Even though there is an existing package that implements adaptive prediction sets (APS) for conformal prediction, coverforest is the first package to implement the Jackknife+-after-Bootstrap (J+ab) method for conformal prediction.
2 Background on Conformal Predictions
Let $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ be training data for a predictive model $\hat{f}$ and $(X_{n+1},Y_{n+1})$ is a new test point, where $X_{i}$’s lie in an input space $\mathcal{X}$ and $Y_{i}$’s lie in an output space $\mathcal{Y}$. Specifically, we let $\mathcal{Y}=\mathbb{R}$ for a regression task and $\mathcal{Y}=\{1,\ldots,C\}$ for a classification task. A prediction set for $Y_{n+1}$ is a set of outputs $\mathcal{C}=\mathcal{C}(X_{i})\subseteq\mathcal{Y}$ that might be input-dependent. For regression tasks we restrict prediction sets to intervals, in which case they are commonly referred to as prediction intervals.
In a conformal prediction task, we specify a miscoverage rate $\alpha\in(0,1)$, with which we aim to construct a prediction set $\mathcal{C}$ that has probability at least $1-\alpha$ to contain the true label:
$P(Y_{n+1}\in\mathcal{C}(X_{n+1}))\geq 1-\alpha.$ (1)
Classical results in conformal prediction literature typically assume that the random variables $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ and $(X_{n+1},Y_{n+1})$ are either i.i.d. or exchangable, that is, their joint distribution are invariant under permutations. Below, we provide a review of well-known conformal prediction methods, and recently proposed conformity that achieve coverage guarantees under one of these assumptions.
Regression
  1. 1.
    Split Conformal [4]. First, we split the data into training and calibration sets, we fit a predictive model $\hat{f}$ on the training data. The method then computes nonconformity scores $R_{i}=|Y_{i}-\hat{f}(X_{i})|$ for each point in the calibration set. For the new test point $X_{n+1}$, the prediction interval is constructed as:
    $\mathcal{C}(X_{n+1})=[\hat{f}(X_{n+1})-q_{\alpha},\hat{f}(X_{n+1})+q_{\alpha}],$
    where $q_{\alpha}$ is the $\lceil(1-\alpha)(n+1)\rceil$-th smallest value of the calibration scores.
  2. 2.
    Cross-Validation+ (CV+) [8]. Split the data into $K$ folds. For each fold $\mathcal{I}_{k}$, we fit a model $\hat{f}^{(-k)}$ on all data except fold $k$ and compute residuals $R_{i}=|Y_{i}-\hat{f}^{(-k)}(X_{i})|$ for all $i\in\mathcal{I}_{k}$. The final prediction interval for a new point is constructed as:
    $\mathcal{C}(X_{n+1})=[q^{-}_{\alpha}\{\hat{f}^{-S(i)}(X_{n+1})-R_{i}\},q^{+}_{\alpha}\{\hat{f}^{-S(i)}(X_{n+1})+R_{i}\}],$
    where $S(i)=k$ if $i\in\mathcal{I}_{k}$, $q^{+}_{\alpha}\{a_{i}\}$ is the $\lceil(1-\alpha)(n+1)\rceil$-th smallest value of $a_{i}$’s and $q^{-}_{\alpha}\{a_{i}\}$ is the $\lfloor\alpha(n+1)\rfloor$-th smallest value of $a_{i}$’s.
  3. 3.
    Jackknife+-after-Bootstrap (J+ab) [6]. This method is similar to CV+ but instead of splitting data into $K$-folds, we sample with or without replacement from the original training set multiple times, and fit a model on each of the subsamples. During prediction, we let $\hat{f}^{(-i)}(X_{i})$ be the aggregated predictions from all models that were fitted on the subsamples that do not contain $(X_{i},Y_{i})$, from which we compute the residuals $R_{i}=|Y_{i}-\hat{f}^{(-i)}(X_{i})|$. The final prediction interval for a new point is then
    $\mathcal{C}(X_{n+1})=[q^{-}_{\alpha}\{\hat{f}^{(-i)}(X_{n+1})-R_{i}\},q^{+}_{\alpha}\{\hat{f}^{(-i)}(X_{n+1})+R_{i}\}],$ (2)
    where $q^{-}_{\alpha}$ and $q^{+}_{\alpha}$ are defined as in CV+. However, this interval with a fixed number of bootstrap samplings $B$ does not achieve a finite-sample coverage guarantee, due to the asymmetry between training and test samples; to see this, observe that each test sample is excluded from exactly $B$ bootstraps, while each training sample is typically excluded from <B<B bootstraps. This issue is resolved by resampling $B$ from a Binomial distribution:
    $\operatorname{Binomial}(\tilde{B},(1-\frac{1}{n+1})^{m}),$ (3)
    where $\tilde{B}$ is a fixed initial value, $n$ is the number of training samples and $m$ is the number of bootstrapped samples. One can show that the interval in (2) with the randomized $B$ restores the symmetry between training and test samples [6].
Classification
  1. 1.
    Split Conformal with APS Score [5]. We split the data into training and calibration sets, then fit a classification model $\hat{f}$ on the training set. An obvious conformity scores are the model’s probability predictions; however, a hypothetical model that assigns 505050-50 probabilities to the most likely and second most likely classes would result in set predictions that only contain the most likely classes, and hence fails to achieve the target coverage level. Alternatively, Adaptive Prediction Sets (APS) [5] has been recently proposed method that is adaptive to data distributions.
    The crux of the APS method is the use of cumulative probability as the conformity score: given the predicted probabilities of a point $(x,y)$ in descending order $\hat{\pi}_{1}(x)\geq\hat{\pi}_{2}(x)\geq\ldots\geq\hat{\pi}_{C}(x)$, the APS score function is the (randomized) generalized inverse quantile:
    $E(x,y,\hat{\pi})=\hat{\pi}_{1}(x)+\ldots+\hat{\pi}_{y-1}(x)+u\cdot\hat{\pi}_{y}(x),\quad u\sim\operatorname{Uniform}(0,1).$ (4)
[…]
    In the case of split conformal prediction, we let $q_{\alpha}$ be the $\lceil(1-\alpha)(n+1)\rceil$-th smallest value of the calibration scores. The prediction set for a new point $X_{n+1}$ is then:
    $\mathcal{C}(X_{n+1})=\left\{1,\ldots,k:k=\inf_{y\in\{1,\ldots,C\}}E(X_{n+1},y,\hat{\pi})\geq q_{\alpha}\right\}.$
    The APS score typically produces larger prediction sets than the probability prediction score; however, the former generally achieves the target coverage level while the latter often fails to do so. In Section 5, we experimentally compare these two scores to confirm this observation.
  2. 2.
    Cross-Validation+ (CV+) with APS Score [5]. The CV+ method for classification aims to improve data efficiency compared to a simple train-calibration split by using a cross-validation scheme. This allows each data point to be used for both training and calibration, but never in the same role simultaneously. The procedure using the Adaptive Prediction Set (APS) score is as follows:
    1. (a)
      Data Partitioning. Split the training data $\{(X_{i},Y_{i})\}_{i=1}^{n}$ into $K$ disjoint folds, denoted by $\mathcal{I}_{1},\ldots,\mathcal{I}_{K}$.
    2. (b)
      Cross-Validation Training. For each fold $k\in\{1,\ldots,K\}$, train a classification model $\hat{\pi}^{(-k)}$ on all data points except those in fold $\mathcal{I}_{k}$. This results in $K$ different models.
    3. (c)
      Leave-one-fold-out Score Calculation. We associate each data point $(X_{i},Y_{i})$, with the model $\hat{\pi}^{(-S(i))}$ where $S(i)=k$ if $i\in\mathcal{I}_{k}$ (i.e., the model that was not trained on the fold containing $i$) to calculate its APS conformity score. Given the predicted probabilities for $(X_{i},Y_{i})$ sorted in descending order $\hat{\pi}_{1}^{(-S(i))}(X_{i})\geq\ldots\geq\hat{\pi}_{C}^{(-S(i))}(X_{i})$, the APS score is $E_{i}=E(X_{i},Y_{i},\hat{\pi}^{-S(i)})$, where the generalized inverse quartile $E$ is defined in (4). This process yields a set of $n$ conformity scores $\{E_{1},\ldots,E_{n}\}$.
    4. (d)
      Prediction Set Construction. For a new test point $X_{n+1}$ and a desired miscoverage rate $\alpha$, we construct the prediction set $\mathcal{C}(X_{n+1})$ as follows: With the conformity scores $E_{1},\ldots,E_{n}$ obtained from the previous step, the $p$-value for a candidate class $y$ is given by:
      $\hat{p}(y)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\left\{E_{i}\geq E(X_{n+1},y,\hat{\pi}^{(-S(i))})\right\}.$ (5)
      The prediction set is then formed by including all classes whose $p$-value is greater than or equal to $\alpha$:
      $\mathcal{C}(X_{n+1})=\{y\in\{1,\ldots,C\}\mid\hat{p}(y)\geq\alpha\}.$ (6)
  3. 3.
    Jackknife+-after-Bootstrap (J+ab) with APS Score [6]. The J+ab method is similar to CV+ but utilizes bootstrapping instead of $K$-fold cross-validation to create leave-one-out models and scores. This approach is particularly well-suited for ensemble methods like random forests.
    1. (a)
      Bootstrap Ensemble. We generate $B$ bootstrap samples from the training data $\{(X_{i},Y_{i})\}_{i=1}^{n}$. For each bootstrap sample $b\in\{1,\ldots,B\}$, train a classification model $\hat{\pi}^{(b)}$.
      As with the J+ab for regression, to ensure a valid finite-sample coverage guarantee, the number of bootstrap samples, $B$, must be randomized according to (3).
    2. (b)
      Leave-one-out Predictions and Scores. For each training point $(X_{i},Y_{i})$, we create an aggregated leave-one-out predictor, $\hat{\pi}^{(-i)}$, by averaging the predictions of all models $\hat{\pi}^{(b)}$ that were trained on bootstrap samples that did not include the point $(X_{i},Y_{i})$. Using this leave-one-out predictor, calculate the APS score $E_{i}=E(X_{i},Y_{i},\hat{\pi}^{(-i)})$, where the generalized inverse quartile $E$ is defined in (4).
    3. (c)
      Prediction Set Construction. For a new test point $X_{n+1}$, the prediction set is constructed by calculating a $p$-value using (5), but with $\hat{\pi}^{(-i)}$ replacing $\hat{\pi}^{(-S(i))}$:
      $\hat{p}(y)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}\left\{E_{i}\geq E(X_{n+1},y,\hat{\pi}^{(-i)})\right\}.$
      The final prediction set is then formed by including all classes with a sufficiently high $p$-value, as in (6):
      $\mathcal{C}(X_{n+1})=\{y\in\{1,\ldots,C\}\mid\hat{p}(y)\geq\alpha\}.$
Regularized Adaptive Prediction Sets (RAPS) [14]. This score was proposed for classification with a lot of classes (say 1000+ classes) where the ordering of the predicted probabilities does not reflect the true order. To this end, the authors propose to add to the APS’s score function a regularization term that penalizes classes with small predicted probabilities (assuming $\hat{\pi}_{1}(x)\geq\hat{\pi}_{2}(x)\geq\ldots\geq\hat{\pi}_{C}(x)$ as before):
$E_{k,\lambda}(x,y)=E(x,y)+\lambda\cdot\max(0,y-k).$ (7)
As with APS, we can use this score function for split conformal, CV+, or J+ab.
Method Theoretical Practical Training Prediction
Coverage Coverage Cost Cost
Regression
Split Conformal $1-\alpha$ $\approx 1-\alpha$ $O(Tdn\log n)$ $O(Tn_{\operatorname{test}}\log n)$
CV+ $1-2\alpha-\epsilon_{K,n}$ $\gtrsim 1-\alpha$ $O(KTdn\log n)$ $O(KTn_{\operatorname{test}}\log n)$
J+ab $1-2\alpha$ $\gtrsim 1-\alpha$ $O(Tdn\log n)$ $O(n*n_{\operatorname{test}})$
Classification
Split + (R)APS $1-\alpha$ $\approx 1-\alpha$ $O(Tdn\log n)$ $O(Tn_{\operatorname{test}}\log n)$
CV+ + (R)APS $1-2\alpha-\epsilon_{K,n}$ $\gtrsim 1-\alpha$ $O(KTdn\log n)$ $O(Cn*n_{\operatorname{test}})$
J+ab + (R)APS $1-2\alpha$ $\gtrsim 1-\alpha$ $O(Tdn\log n)$ $O(Cn*n_{\operatorname{test}})$
Table 1: Comparison of coverage guarantees and computational costs for different conformal random forest methods. Here, the notation $\gtrsim 1-\alpha$ means “sometimes worse than $1-\alpha$”, $n$ is the number of training samples, $n_{\operatorname{test}}$ is the number of test samples, $T$ is the number of trees, $C$ is the number of classes, $K$ is the number of folds in CV+, and $\epsilon_{K,n}=\min\left\{2(1-1/K)/(n/K+1),(1-K/n)/(K+1)\right\}$. We assume that $T\ll n$ and $C\leq T$.
In Table 1, we compare coverage guarantees and computational costs across different conformal prediction methods for random forests. While CV+ and J+ab theoretically guarantee coverages of $\approx 1-2\alpha$, they typically achieve $\geq 1-\alpha$ coverage in practice. One downside of CV+ and J+ab with the APS and RAPS scores is the notably long prediction times due to the pairwise comparisons between train and test conformity scores across all classes, which contribute to $Cn*n_{\text{test}}$ computational cost.
2.1 Related work
The coverforest package implements conformal predictions [3, 4, 15, 1, 2] using random forest as the underlying model [7]. For similar work, conformity scores of bootstrapped regression based on iterative reweighted least squares is briefly discussed in [4]. For random forest classification, two conformity scores based on out-of-bag predictions were proposed in [16]. However, a finite-sample coverage guarantee cannot be obtained as they do not preserve exchangability [6].
Nonconformity scores based on out-of-bag (OOB) predictions were introduced in [17]. The experiments revealed that these scores yield smaller prediction intervals compared to neural networks and $k$-nearest neighbors. Later, it is confirmed in [18] that these intervals provide asymptotic coverage guarantee, which is only attained with infinitely many data. Further works explored various aspects of out-of-bag calibration for conformal predictions [19, 20, 21]. coverforest’s implementation of Jackknife+-after-bootstrap uses a randomized number of bootstraps in order to achieve a finite-sample coverage guarantee, which is proved in [6].
A conformal regression method related to CV+ is Jackknife+ [8] where the regression model is fitted on leave-one-out datasets. It is a special case of CV+ where each fold only consists of a single training point, and the number of folds equals the number of training samples. coverforest can be used with the Jackknife+ method by specifying method="cv" and cv equals the number of training samples. Another method is the conformalized quantile regression (CQR) which constructs prediction intervals based on classical quantile regression. Implementing CQR is one of our future plans where we can incorporate existing implementation of conformal quantile random forest [22] such as quantile-forest [23].
For classification, given a model that outputs probability estimates $\hat{\pi}_{j}(x)\approx P(Y=j|X=x)$, a basic conformal prediction method can be devised based on a score function $s(x,y)=1-\hat{\pi}_{y}(x)$. The prediction set for a new point $X_{n+1}$ is then given by $\mathcal{C}(X_{n+1})=\left\{y:s(X_{n+1},y)\leq q_{\alpha}\right\}$, where $q_{\alpha}$ is the $\lceil(1-\alpha)(n+1)\rceil$-th smallest value of the calibration scores [24, 4]. However, this method tend to outputs an empty set when the model is uncertain about the prediction, for example, $q_{\alpha}=0.2$ and the predicted probabilities of two classes are 0.5 each. This issue can be remedied by instead using the cumulative probability as in the APS method [5]. Our future plans include implementing multilabel classification with finite-sample coverage guarantee via conformal risk control [25, 26, 27].
There are several software packages that implement various types of conformal predictions, such as inductive conformal prediction and mondrian conformal prediction. However, coverforest is the first package to implement CV+, J+ab and split conformal with both APS and RAPS scores for set predictions tasks. See Table 1 for a feature comparison between coverforest and other conformal prediction packages. We will also see in Section 5 that coverforest’s implementations of APS and RAPS-based set predictions are faster than the alternative (MAPIE). by 2 to 5 times
Score Method
Package APS RAPS CV+ J+ab Split Inductive Mondrian
nonconformist [28]
Orange3-Conformal [29]
conforest [17]
crepes [30]
MAPIE [31]
coverforest
Table 2: Feature comparison between various conformal prediction software packages for set prediction tasks.
3 Implementation
3.1 API description
Figure 1: Class structure and the metrics in coverforest.
coverforest’s provides two main model classes with familiar scikit-learn-style API, with the standard fit() and predict() methods, as shown in Figure 1. The most important parameters of both classes are:
  • n_estimators: The number of sub-estimators in the forest
  • method: The conformal prediction methods, which can be one of cv (CV+), bootstrap (Jackknife+-after-bootstrap), and split, (split conformal)
  • cv: The number of folds in the CV+ method
There are also several parameters that are only specific to APS and RAPS:
  • k_init_ and lambda_init_: The regularization parameters in RAPS. The APS score is equivalent to setting k_init_=0 and lambda_init_=0. Sutomatic searching using the method described in [14] is also available by setting k_init_="auto" and lambda_init_="auto"
  • allow_empty_set: whether to allow empty set predictions
  • randomized: whether to apply the APS’s randomization scheme to include an additional class in prediction sets as described in Section 2.
There are only one parameter that is specific to the Jackknife+-after-bootstrap method:
  • resample_n_estimators: whether to draw the number of trees from the Binomial distribution (3), in which case n_estimators is taken as the first parameter $\tilde{B}$ of the distribution.
We now go over the methods in both model classes.
The fit() method fits each tree on the training data. The trees are fitted in parallel via a helper class joblib.Parallel() [12]. The fitted trees are then used to calculate for each training sample the out-of-box (OOB) predictions, which are then used to calculate the calibration scores. During fitting, the method maintains $\mathcal{T}_{i}$ for each sample $(X_{i},Y_{i})$, where $\mathcal{T}_{i}$ is the collection of trees not fitted on this particular sample.
The predict() method first queries predictions from all trees for any given test sample $(X_{n+1},Y_{n+1})$, where the queries are performed in parallel via joblib.Parallel(). Subsequently, for each training instance $i$, it aggregates predictions exclusively from trees in $\mathcal{T}_{i}$, converting these into a conformity score. This score is then compared with the calibration scores of $(X_{i},Y_{i})$ to produce the final predictions.
For moderately large training and test sets, calculating the conformity scores across all pairs of training and test samples can be quite expensive. We contribute to performance improvement by implementing cross-calculations in Cython with parallel processing enabled. Specifically, the _giqs.pyx script in the package contains optimized code that speeds up calculation of cross-conformity scores, resulting in substantially faster prediction time than the previous package. Users can specify the number of threads through the num_threads parameter in the predict() method.
The CoverForestClassifier includes the search_k_and_lambda() method for automatic parameter optimization. This method determines optimal values for $k$ and $\lambda$ according to the approach detailed in [14].
coverforest also has a metric module that provides convenient functions for evaluating the prediction sets/intervals by measuring their average size and empirical coverage (e.g. the proportion of prediction sets/intervals that contains the true class/value).
3.2 Computational complexity and memory usage
As summarized in Table 1, both CV+ and J+ab methods require $Cn*n_{\text{test}}$ memory and computations for the calculation of conformity scores. To reduce the memory, the user could split the test set into several chunks and make predictions on each chunk. Alternatively, the user could use the split conformal method which requires only $n*n_{\text{classes}}$ memory and $Tn_{\text{test}}\log n$ computations while being able to achieve satisfactorily small set sizes on a large dataset.
3.3 Source code
The source code for the coverforest is hosted on GitHub at https://github.com/donlap/coverforest. The code follows the PEP8 guidelines, and all docstrings of publicly exposed classes comply with the NumPy docstring style. Both CoverForestClassifier and CoverForestRegressor pass the comprehensive unit tests provided by scikit-learn. The documentation includes installation guides, user manual, and API reference, generated using sphinx and available online at https://donlapark.github.io/coverforest. The package is BSD-licensed and requires only scikit-learn version 1.6.1 or greater.
4 Examples
We demonstrate how to use coverforest on a toy dataset for set predictions.
Conformal predictions with CV+ and RAPS score
We train a conformal random forest using CV+ with the RAPS score. To use the CV+ method, initialize a CoverForestClassifier model with method="cv". Here, we set cv=5 to split the data into five folds. As a result, five random forests will be fitted on the five combinations of 4-folds. We also specify that each forest has 200 trees by setting n_estimators=200.
from coverforest.metrics import average_set_size_loss
from coverforest.metrics import classification_coverage_score
from sklearn.datasets import load_digits
from sklearn.model_selection import train_test_split
X, y = load_digits(return_X_y=True)
X_train, X_test, y_train, y_test = train_test_split(X, y, test_size=0.2)
clf = CoverForestClassifier(n_estimators=200, method="cv", cv=5)
Training and making predictions
We train the random forest by calling the fit() method, and make a prediction by calling the predict() method. During prediction, we choose the miscoverage rate at $0.05$ by specifying alpha=0.05. The output of the predict() method is a tuple: The first item contains the single class predictions by the random forest, and the second item contains the set predictions.
y_pred, y_sets = clf.predict(X_test, alpha=0.05)
# y_pred[:5] = [2, 8, 2, 6, 6]
# y_sets[:5] = [array([2]), array([2, 5, 8]), array([2]),
# array([5, 6]), array([6])]
Computing average set size and coverage
We may use the provided metrics functions to evaluate the set predictions against the true classes.
coverage = classification_coverage_score(y_test, y_sets)
# avg_set_size = 1.589
# coverage = 0.964
The small average set size indicates that the model can make predictions with high confidence. The coverage tells us that $96.4\%$ of the prediction sets contain the true classes, which is above our specified level of $95\%$.
Conformal predictions using Jackknife+-after-bootstrap with the APS score
In the second example, we train a conformal random forest using J+ab with the APS score. To use J+ab, initialize a CoverForestClassifier model with method="bootstrap". To use the APS method, set k_init=0 and lambda_init=0. In contrast to CV+ which has multiple forests, J+ab has only a single forest. Here, we specify that this forest has 400 trees.
n_estimators=400, method="bootstrap", k_init=0, lambda_init=0
clf.fit(X_train, y_train)
y_pred, y_sets = clf.predict(X_test, alpha=0.05)
avg_set_size = average_set_size_loss(y_test, y_sets)
coverage = classification_coverage_score(y_test, y_sets)
# y_pred[:5] = [2, 8, 2, 6, 6]
# y_sets[:5] = [array([2]), array([2, 5, 8, 9]),
# array([2]), array([6]), array([6])]
# avg_set_size = 1.761
# coverage = 0.968
Compared to the CV+ method, J+ab produces prediction sets with a larger average with practically the same coverage.
5 Experiments and Discussions
To validate our implementation, we benchmark coverforest’s three methods: CV+, J+ab, and split conformal on conformal classfication and regression tasks. We compare our implementations against the most recent packages for conformal predictions, namely the crepes and MAPIE packages.
5.1 Classification experiment
We evaluate the predictions and run times of CoverForestClassifier on five classification datasets, namely Mice [32], WineQuality [33], Bean [34], MNIST, and Helena [35] datasets. See Table 3 for the datasets’ characteristics.
We aim to evaluate the methods’ abilities to achieve the desired level of coverage even with a few training points. To this end, we randomly select 20% of the dataset as the test set (except the MNIST dataset whose test set has already been prepared), and 200 rows of the remaining data as the training set. For the split conformal method, we reserve 100 rows of the training set as the calibration set. We then fit the models on the training set with the target miscoverage level $\alpha\in\{0.05,0.1,0.2\}$. We repeat the experiment 50 times for each method and compare the empirical coverage (the proportion of prediction sets that contain the true class), and average set size.
For CV+, we split the training data into 10 folds, and each random forest has 30 trees. For J+ab and split conformal, the random forest has 100 trees.
Experiment 1 & 2 Experiment 3
Classes Features Train Test Train Test
Mice [32] 8 80 200 216 861 216
WineQuality [33] 7 11 200 400 3918 980
Bean [34] 7 16 200 2723 5000 1250
MNIST [36] 10 784 200 10000 5000 1250
Helena [35] 100 27 1033 300 904 300
Table 3: The datasets used in the conformal classification experiments. The smaller training set of the Helena dataset in Experiment 3 is due to the memory limit of the MAPIE package.
We also compare our methods against crepes [37], which uses the classifier’s predicted probabilities as the conformity scores; however, such scores are not adaptive in the sense that they tend to result in coverages that are over or under the nominal level, depending on the accuracy of the classifier.
crepes offers the OOB and split conformal methods. In our experiment, we run both methods with a random forest with 100 trees as the base model.
As the MAPIE library implements the same CV+ and split conformal methods as ours, we do not run MAPIE in this experiment. Nonetheless, we will compare our run times against MAPIE in a later experiment.
For Experiment 1 and 2, we also perform hypothesis testing to verify that the empirical coverage probabilities exceed the nominal coverage levels, and to compare the average set sizes between any pair of the methods. The results and discussion of the tests can be found in A.
Figure 2: Coverage probabilities of CoverForestClassifier’s set predictions on the test set, repeated 50 times for each of the CV+, J+ab and split conformal methods with the APS and RAPS scores. The results of crepes’s methods are also included.
Experiment 1. Empirical coverage
Figure 2 presents the empirical coverages achieved using both the APS (4) and RAPS (7) scores. The plots consistently demonstrate that CV+ provides the largest coverage. J+ab consistently yields slightly lower coverages than CV+. On the Mice, Bean, and Helena datasets, the OOB method from crepes results in notably large coverage across all values of $\alpha$. In contrast, the split conformal method provides the lowest and most inconsistent coverages on these same datasets.
We now turn our attention to the effect of regularization on the empirical coverages. In this regard, the CV+ and J+ab methods, when paired with the RAPS score, yield slightly reduced coverages for $\alpha=0.1$ and $0.2$. This difference becomes more pronounced in datasets with a large number of classes, as is evident in the results for the MNIST and Helena datasets.
Finally, we assess whether these methods achieve their nominal coverage levels. The plots indicate that the median coverages for both the CV+ and J+ab methods exceed these target levels in the majority of cases. This outcome demonstrates the data efficiency of these two approaches, which allows them to successfully reach the desired coverage.
Experiment 2. Average set size
Figure 3 illustrates the average prediction set sizes produced by each method. Across all datasets, we observe a general trend where CV+ produces the largest average set sizes, followed by J+ab, and then the split conformal method. The split and OOB methods from crepes yield the smallest set sizes on the Mice, Bean, and MNIST datasets. It is important to recall from the previous experiment, however, that these methods did not achieve the nominal coverage levels on the MNIST dataset.
We now examine the impact of regularization on the average set sizes. For all datasets, the application of the RAPS score contributes to a reduction in the set sizes for the CV+, J+ab, and split conformal methods. This decrease in set size is substantially more noticeable for datasets with a large number of classes, as demonstrated by the results on the Helena dataset.
Figure 3: Average size of CoverForestClassifier’s set predictions on the test set, repeated 50 times for each of the CV+, J+ab and split conformal methods with the APS and RAPS scores. The results of crepes’s methods are also included.
Experiment 3. Run time comparison with MAPIE
In this experiment, we compare the run time of CoverForestClassifier against MAPIE [31]. As far as we know, it is currently the only package that provides the CV+ methods with the APS and RAPS scores (the J+ab method is also provided in MAPIE but only for regression tasks).
We do not compare our methods against the crepes library here as it does not provide the CV+ and J+ab methods, nor the APS and the RAPS scores.
For coverforest and MAPIE, we train a random forest and make conformal predictions using CV+, J+ab and split conformal methods, all with the APS score. The numbers of training and test instances in this experiment are indicated in Table 3. We run CV+ with 100 forests with 10 trees each, and J+ab and split conformal with 900 trees. And we fix $\alpha=0.05$. We run each method for 30 repetitions each and report the average training and prediction times.
Figure 4 compares the run times of both packages (note that at the time of this study, MAPIE had not implemented J+ab with the APS score). It shows that coverforest consistently runs faster than MAPIE in both training and prediction, especially with large datasets like WineQuality and MNIST. The improvements are even more pronounced on datasets with many classes. Specifically, on the Helena dataset, coverforest training time is 14 times faster for the CV+ method, 8 times faster for the split method; and the testing time is 72 times faster for the CV+ method, and 30 times faster for the split method. These results indicate the performance advantage to our optimized Numpy code during the training phase, and the optimized Cython code during the prediction phase.
Figure 4: Training and prediction times of coverforest’s and MAPIE’s implementations of the CV+, J+ab and split conformal methods with the APS score and $\alpha=0.05$. Note that MAPIE does not implement the J+ab method for classification.
5.2 Regression experiment
Experiment 1 & 2 Experiment 3
Ranges Features Train Test Train Test
Housing [38] [0.15, 5.0] 8 200 1000 16512 4128
Concrete [39] [2.33, 82.6] 8 200 206 824 206
Bike [40] [1, 638] 12 200 1000 13903 3476
Crime [41] [0.0, 1.0] 122 200 399 1595 399
Table 4: The datasets used in our experiment.
We now evaluate CoverForestRegressor for interval predictions on four regression datasets. The characteristics of the benchmark datasets are detailed in Table 4.
For CV+, we split the training data into 10 folds, and each random forest has 30 trees. For J+ab and split conformal, the random forest has 100 trees. We randomly sample 20% of each dataset for the test set and 200 instances of the remaining data for the training set. For each dataset, we run each method for 50 repetitions and calculate the coverage and average interval length. The numbers of training and test instances in this experiment are indicated in Table 4. We fix $\alpha=0.05$. For each method, we perform training and prediction for 30 repetitions and compare the average run times to those of crepes and MAPIE.
For Experiment 1 and 2, we also perform hypothesis testing to verify that the empirical coverage probabilities exceed the nominal coverage levels, and to compare the interval lengths between any pair of the methods. The results and discussion of the tests can be found in A.
Experiment 1 & 2. Empirical coverage and average interval length
Figure 5: Coverage and average length of CoverForestRegressor’s interval predictions on the test set, repeated 50 times for each of the CV+, J+ab and split conformal methods with the residuals as the conformity scores. Results of crepes’s J+ab method are also included.
Figure 5 shows the coverages and average interval lengths of the three methods. Across all datasets, both CV+ and J+ab produce prediction intervals with small variances compared to the split conformal, with J+ab consistently producing intervals of smaller average lengths and smaller coverages than CV+. Even though both crepes’s methods provide coverages that achieve the nominal levels, their interval lengths are substantially longer than those from the other methods.
Experiment 3. Run time comparison with crepes and MAPIE
Figure 6 presents a comparison of the run times for coverforest, crepes, and MAPIE on conformal regression tasks. The results clearly show that coverforest is substantially faster than the other packages during the training phase across all evaluated methods.
Regarding prediction time, crepes is generally the fastest package, noticeably with the bootstrap (J+ab) method; this is due to the implementation of the Mondrian conformal method, which only requires a single prediction for each test point. In contrast, the methods implemented in our package and in MAPIE necessitate making multiple predictions for each test point (more precisely, for each training point, its out-of-bag trees are used to make predictions for the test point). Nonetheless, it is important to note that, unlike in classification, prediction in conformal regression takes substantially less time than training.
Figure 6: Training and prediction times of coverforest, crepes and MAPIE’s implementations of the CV+, J+ab and split conformal methods with the residuals as the conformity scores and $\alpha=0.05$. Note that crepes does not implement the CV+ method.
6 Conclusion
In this work, we introduce coverforest, a lightweight package that bridges the gap between scikit-learn’s optimized random forest implementation and modern conformal prediction techniques. By leveraging both cross-validation+ and Jackknife+-after-bootstrap approaches, coverforest enables uncertainty quantification through distribution-free prediction intervals and prediction sets.
Our empirical evaluation demonstrates that coverforest not only achieves the desired coverage guarantees but is also faster than existing implementation of conformal predictions with random forests. Given random forest’s established effectiveness on tabular data, coverforest provides a practical solution for deployment of conformal prediction in real-world applications. Future work could explore extensions to other ensemble methods and conformal prediction techniques while maintaining the package’s emphasis on simplicity and efficiency.
Finally, we acknowledge a current limitation of our package regarding computational scalability. The implementation is highly optimized for multi-core CPU execution via Cython, but it does not currently support GPU acceleration or distributed computing. A significant avenue for future work is to extend the framework to leverage GPU-native libraries, such as cuML [42], or distributed platforms like Dask [43] and Ray [44] While this would require a substantial change to the codebase, it would greatly enhance the applicability of our methods to large-scale datasets that exceed the capabilities of a single machine.
Acknowledgment
This work was supported by Chiang Mai University, Thailand.
A Significant testing of the experiments’ results
To verify the coverage guarantees and compare the methods in terms of the average set sizes, we perform two statistical tests on the results in Section 5:
  1. 1.
    One-sided $t$-test to verify that the empirical coverage probability for each combination of method, dataset and $\alpha$ exceeds the target of $1-\alpha$.
  2. 2.
    Two-sample $t$-test to compare the average set sizes/average interval lengths for each pair of methods in each dataset.
For the one-sided test, we also provide $p$-values in hypotheses. For both tests, we use the star symbols: “*”, “**” and “***” to indicate that the test’s conclusion is significant at level $0.05,0.01$ and $0.001$, respectively. For the two-sample $t$-test, the star symbol indicates the smallest significance level that holds for all values of nominal coverage $\alpha$.
A.1 Classification
The results of the one-sided tests for the five classification datasets are shown in Table 5, 6, 7, 8 and 9. Those of the two-sample tests are shown in Table 10, 11, 12, 13 and 14. Here are some observations from these results:
  • While the split and crepes’ methods often produce smaller sets on average than the CV+ and J+ab methods, they frequently fail to achieve the target coverage levels, which defeats the purpose of conformal prediction.
  • For the CV+ and J+ab methods, using RAPS scores as opposed to APS scores leads to a significant reduction in prediction set sizes, particularly for the MNIST and Helena datasets which have large numbers of classes.
  • In some cases, J+ab produces smaller set on average than CV+, but not the other way around.
Table 5: One-Sided $t$-test for Coverage > $1-\alpha$ for the Mice dataset
$\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$
Model
CV+ (APS) 0.970 (0.000)*** 0.925 (0.000)*** 0.832 (0.000)***
CV+ (RAPS) 0.970 (0.000)*** 0.925 (0.000)*** 0.832 (0.000)***
J+ab (APS) 0.961 (0.000)*** 0.920 (0.000)*** 0.826 (0.000)***
J+ab (RAPS) 0.961 (0.000)*** 0.920 (0.000)*** 0.825 (0.000)***
Split (APS) 0.953 (0.204) 0.907 (0.081) 0.814 (0.021)*
Split (RAPS) 0.950 (0.532) 0.907 (0.076) 0.815 (0.020)*
crepes (split) 0.970 (0.000)*** 0.898 (0.669) 0.737 (1.000)
crepes (oob) 0.984 (0.000)*** 0.940 (0.000)*** 0.826 (0.000)***
Table 6: One-Sided $t$-test for Coverage > $1-\alpha$ for the WineQuality dataset
$\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$
Model
CV+ (APS) 0.961 (0.000)*** 0.921 (0.000)*** 0.848 (0.000)***
CV+ (RAPS) 0.957 (0.000)*** 0.917 (0.000)*** 0.842 (0.000)***
J+ab (APS) 0.951 (0.218) 0.912 (0.000)*** 0.830 (0.000)***
J+ab (RAPS) 0.948 (0.965) 0.909 (0.000)*** 0.821 (0.000)***
Split (APS) 0.909 (1.000) 0.800 (1.000) 0.638 (1.000)
Split (RAPS) 0.909 (1.000) 0.800 (1.000) 0.629 (1.000)
crepes (split) 0.956 (0.000)*** 0.892 (0.893) 0.621 (1.000)
crepes (oob) 0.946 (0.998) 0.909 (0.029)* 0.722 (1.000)
Table 7: One-Sided $t$-test for Coverage > $1-\alpha$ for the Bean dataset
$\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$
Model
CV+ (APS) 0.972 (0.000)*** 0.934 (0.000)*** 0.842 (0.000)***
CV+ (RAPS) 0.971 (0.000)*** 0.933 (0.000)*** 0.834 (0.000)***
J+ab (APS) 0.967 (0.000)*** 0.929 (0.000)*** 0.838 (0.000)***
J+ab (RAPS) 0.966 (0.000)*** 0.927 (0.000)*** 0.824 (0.000)***
Split (APS) 0.922 (1.000) 0.875 (1.000) 0.783 (1.000)
Split (RAPS) 0.917 (1.000) 0.872 (1.000) 0.780 (1.000)
crepes (split) 0.951 (0.256) 0.913 (0.000)*** 0.694 (1.000)
crepes (oob) 0.974 (0.000)*** 0.920 (0.000)*** 0.831 (0.000)***
Table 8: One-Sided $t$-test for Coverage > $1-\alpha$ for the MNIST dataset
$\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$
Model
CV+ (APS) 0.971 (0.000)*** 0.925 (0.000)*** 0.824 (0.000)***
CV+ (RAPS) 0.970 (0.000)*** 0.921 (0.000)*** 0.803 (0.069)
J+ab (APS) 0.958 (0.000)*** 0.909 (0.000)*** 0.808 (0.002)**
J+ab (RAPS) 0.956 (0.000)*** 0.905 (0.005)** 0.786 (1.000)
Split (APS) 0.927 (1.000) 0.851 (1.000) 0.750 (1.000)
Split (RAPS) 0.927 (1.000) 0.851 (1.000) 0.743 (1.000)
crepes (split) 0.935 (1.000) 0.869 (1.000) 0.684 (1.000)
crepes (oob) 0.948 (0.823) 0.865 (1.000) 0.682 (1.000)
Table 9: One-Sided $t$-test for Coverage > $1-\alpha$ for the Helena dataset
$\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$
Model
CV+ (APS) 1.000 (0.000)*** 0.973 (0.000)*** 0.858 (0.000)***
CV+ (RAPS) 0.997 (0.000)*** 0.953 (0.000)*** 0.837 (0.000)***
J+ab (APS) 0.998 (0.000)*** 0.947 (0.000)*** 0.848 (0.000)***
J+ab (RAPS) 0.984 (0.000)*** 0.937 (0.000)*** 0.832 (0.000)***
Split (APS) 0.983 (0.000)*** 0.914 (0.000)*** 0.812 (0.000)***
Split (RAPS) 0.983 (0.000)*** 0.910 (0.000)*** 0.806 (0.006)**
crepes (split) 0.971 (0.000)*** 0.938 (0.000)*** 0.855 (0.000)***
crepes (oob) 0.988 (0.000)*** 0.976 (0.000)*** 0.878 (0.000)***
Table 10: Pairwise $t$-test for Average Set Size for the Mice dataset
CV+ (A) CV+ (R) J+ab (A) J+ab (R) Spl (A) Spl (R) Cr (spl) Cr (oob)
CV+ (A) - - $\blacktriangle$ ** $\blacktriangle$ ** $\blacktriangle$ ** $\blacktriangle$ *** $\blacktriangle$ ***
CV+ (R) - - - $\blacktriangle$ ** $\blacktriangle$ * $\blacktriangle$ *** $\blacktriangle$ ***
J+ab (A) - - $\blacktriangle$ * $\blacktriangle$ ** $\blacktriangle$ ** $\blacktriangle$ *** $\blacktriangle$ ***
[…]
Spl (A) $blacktriangledown ** \blacktriangledown ** - - \blacktriangle *** \blacktriangle ***$
Spl (R) $blacktriangledown * \blacktriangledown ** - - \blacktriangle *** \blacktriangle ***$
Cr (spl) $blacktriangledown *** \blacktriangledown *** \blacktriangledown *** \blacktriangledown *** \blacktriangledown *** \blacktriangledown ***$
Cr (oob) $blacktriangledown *** \blacktriangledown *** \blacktriangledown *** \blacktriangledown *** \blacktriangledown *** \blacktriangle ***$
Table 11: Pairwise $t$-test for Average Set Size for the WineQuality dataset
CV+ (A) CV+ (R) J+ab (A) J+ab (R) Spl (A) Spl (R) Cr (spl) Cr (oob)
[…]
J+ab (R) $blacktriangledown *** \blacktriangledown * \blacktriangle *** \blacktriangle *** \blacktriangledown *** \blacktriangledown ***$
Spl (A) $blacktriangledown *** \blacktriangledown *** \blacktriangledown *** - \blacktriangledown *** \blacktriangledown ***$
Spl (R) $blacktriangledown *** \blacktriangledown *** \blacktriangledown *** - \blacktriangledown *** \blacktriangledown ***$
Cr (spl) - $\blacktriangle$ ** $\blacktriangle$ ** $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** -
Cr (oob) $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** -
Table 12: Pairwise $t$-test for Average Set Size for the Bean dataset
CV+ (A) CV+ (R) J+ab (A) J+ab (R) Spl (A) Spl (R) Cr (spl) Cr (oob)
CV+ (A) - - $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ ***
CV+ (R) - - $\blacktriangle$ * $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ ***
J+ab (A) - - - $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ ***
[…]
Spl (A) $blacktriangledown *** \blacktriangledown *** \blacktriangledown *** - \blacktriangle *** \blacktriangle *$
Spl (R) $blacktriangledown *** \blacktriangledown *** \blacktriangledown *** - \blacktriangle *** -$
Cr (spl) $blacktriangledown *** \blacktriangledown *** \blacktriangledown *** \blacktriangledown *** \blacktriangledown *** -$
Cr (oob) $blacktriangledown *** \blacktriangledown *** \blacktriangledown *** \blacktriangledown * - -$
Table 13: Pairwise $t$-test for Average Set Size for the MNIST dataset
CV+ (A) CV+ (R) J+ab (A) J+ab (R) Spl (A) Spl (R) Cr (spl) Cr (oob)
CV+ (A) - $\blacktriangle$ * $\blacktriangle$ *** - - $\blacktriangle$ *** $\blacktriangle$ ***
CV+ (R) - - $\blacktriangle$ *** - - $\blacktriangle$ *** $\blacktriangle$ ***
[…]
Spl (A) - - - - - $\blacktriangle$ *** $\blacktriangle$ ***
Spl (R) - - - - - $\blacktriangle$ *** $\blacktriangle$ ***
Cr (spl) $blacktriangledown *** \blacktriangledown *** \blacktriangledown *** \blacktriangledown *** \blacktriangledown *** -$
Cr (oob) $blacktriangledown *** \blacktriangledown *** \blacktriangledown *** \blacktriangledown *** \blacktriangledown *** -$
Table 14: Pairwise $t$-test for Average Set Size for the Helena dataset
CV+ (A) CV+ (R) J+ab (A) J+ab (R) Spl (A) Spl (R) Cr (spl) Cr (oob)
CV+ (A) $\blacktriangle$ *** - $\blacktriangle$ *** - - - -
[…]
J+ab (A) - - $\blacktriangle$ *** - - - -
J+ab (R) $blacktriangledown *** \blacktriangledown * - - \blacktriangledown ***$
Spl (A) - - - $\blacktriangle$ * - - -
Spl (R) - - - - - - -
[…]
Cr (oob) - - - $\blacktriangle$ *** - - $\blacktriangle$ ***
A.2 Regression
The results of the one-sided tests for the four regression datasets are shown in Table 15, 16, 17 and 18. Those of the two-sample tests are shown in Table 19, 20, 21 and 22. Here are some observations from these results:
  • On the Concrete and Bike datasets, all methods except crepes (split) achieve the target coverages at $0.001$ significance level. However, for the Housing dataset, only CV+ achieves the target coverage at $0.05$ significance level or less.
  • In all cases, both crepes’ methods produce larger intervals on average than the CV+, J+ab and split methods.
  • In all cases, the J+ab method produces smaller intervals on average than the split method.
  • On all but the Crime dataset, the J+ab method produces smaller intervals on average than the CV+ method.
Table 15: One-Sided $t$-test for Coverage > $1-\alpha$ for the Housing dataset
$\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$
Model
CV+ 0.951 (0.039)* 0.910 (0.000)*** 0.823 (0.000)***
J+ab 0.944 (1.000) 0.897 (0.996) 0.805 (0.000)***
Split 0.951 (0.283) 0.896 (0.867) 0.801 (0.432)
crepes (split) 0.949 (1.000) 0.898 (1.000) 0.794 (1.000)
crepes (oob) 0.953 (0.000)*** 0.900 (0.196) 0.798 (1.000)
Table 16: One-Sided $t$-test for Coverage > $1-\alpha$ for the Concrete dataset
$\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$
Model
CV+ 0.977 (0.000)*** 0.941 (0.000)*** 0.877 (0.000)***
J+ab 0.969 (0.000)*** 0.932 (0.000)*** 0.855 (0.000)***
Split 0.982 (0.000)*** 0.944 (0.000)*** 0.857 (0.000)***
crepes (split) 0.941 (1.000) 0.907 (0.000)*** 0.809 (0.000)***
crepes (oob) 0.963 (0.000)*** 0.915 (0.000)*** 0.817 (0.000)***
Table 17: One-Sided $t$-test for Coverage > $1-\alpha$ for the Bike dataset
$\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$
Model
CV+ 0.980 (0.000)*** 0.953 (0.000)*** 0.879 (0.000)***
J+ab 0.979 (0.000)*** 0.944 (0.000)*** 0.861 (0.000)***
Split 0.973 (0.000)*** 0.935 (0.000)*** 0.850 (0.000)***
crepes (split) 0.953 (0.000)*** 0.900 (0.074) 0.806 (0.000)***
crepes (oob) 0.953 (0.000)*** 0.909 (0.000)*** 0.811 (0.000)***
Table 18: One-Sided $t$-test for Coverage > $1-\alpha$ for the Crime dataset
$\alpha=0.05$ $\alpha=0.1$ $\alpha=0.2$
Model
CV+ 0.961 (0.000)*** 0.896 (0.993) 0.797 (0.949)
J+ab 0.958 (0.000)*** 0.888 (1.000) 0.785 (1.000)
Split 0.963 (0.000)*** 0.909 (0.025)* 0.792 (0.980)
crepes (split) 0.953 (0.000)*** 0.885 (1.000) 0.788 (1.000)
crepes (oob) 0.963 (0.000)*** 0.909 (0.000)*** 0.785 (1.000)
Table 19: Pairwise $t$-test for Average Interval Length for the Housing dataset
CV+ J+ab Split crepes (split) crepes (oob)
CV+ $\blacktriangle$ *** $blacktriangledown *** \blacktriangledown ***$
J+ab $blacktriangledown *** \blacktriangledown *** \blacktriangledown ***$
[…]
crepes (split) $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** -
crepes (oob) $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** -
Table 20: Pairwise $t$-test for Average Interval Length for the Concrete dataset
CV+ J+ab Split crepes (split) crepes (oob)
CV+ $\blacktriangle$ *** $blacktriangledown *** \blacktriangledown ***$
J+ab $blacktriangledown *** \blacktriangledown *** \blacktriangledown ***$
[…]
crepes (split) $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** -
crepes (oob) $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** -
Table 21: Pairwise $t$-test for Average Interval Length for the Bike dataset
CV+ J+ab Split crepes (split) crepes (oob)
[…]
J+ab $blacktriangledown * \blacktriangledown *** \blacktriangledown ***$
[…]
crepes (oob) $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ ***
Table 22: Pairwise $t$-test for Average Interval Length for the Crime dataset
CV+ J+ab Split crepes (split) crepes (oob)
[…]
J+ab - $blacktriangledown *** \blacktriangledown ***$
[…]
crepes (split) $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** -
crepes (oob) $\blacktriangle$ *** $\blacktriangle$ *** $\blacktriangle$ *** -
B Cython Implementation for Cross-Conformity Scores
This appendix details the Cython implementations responsible for the computational speed-up of coverforest. The performance gains are primarily achieved through parallel processing with OpenMP and optimized C++ memory management that avoids Python’s overhead.
We present two Cython functions that calculate the cross-conformity scores:
  1. 1.
    _compute_predictions_split() for the split-conformal method
  2. 2.
    _compute_test_giqs_cv() for the CV+ and J+ab methods.
For _compute_predictions_split() the steps to calculate the cross-conformity scores for each test sample are:
  1. 1.
    Sort Probabilities: The class probabilities $\hat{\pi}$ are sorted in descending order to establish their ranks.
  2. 2.
    Compute Cumulative Sum: The cumulative sum of the sorted probabilities is calculated.
  3. 3.
    Apply RAPS Regularization: The regularization penalty, controlled by lambda_star and k_star, is added to the cumulative scores. This penalty increases for each class included beyond the k_star-th class, discouraging overly large prediction sets.
  4. 4.
    Determine Set Size (L): The function finds the smallest $L$ such that the RAPS score for the $L$-th ranked class is greater than the threshold $\tau$.
  5. 5.
    Randomization: If randomized is true, a random draw $u$ is used to decide whether to include the $L$-th class. This is crucial for achieving exact marginal coverage by tightening the prediction sets.
The code for _compute_predictions_split() is shown below. The most crucial part is the with nogil , parallel() that allows us to release the global interpreter lock (GIL) and calculate the scores in parallel:
float64_t[:,::1] out,
float64_t tau,
int64_t k_star,
float32_t lambda_star,
bint randomized,
bint allow_empty_sets,
intp_t num_threads,
intp_t random_state):
"""Construct prediction sets from calibration set’s generalized inverse quantile comformity scores (giqs).
# Variable declarations
cdef Py_ssize_t i, j
cdef Py_ssize_t n_samples = oob_pred.shape[0]
cdef Py_ssize_t n_classes = oob_pred.shape[1]
cdef int32_t* indices
cdef float64_t* sorted_scores, *cumsum
cdef float64_t* *penalties, *penalties_cumsum
cdef int32_t* L
cdef float32_t* U
# Release the GIL and start parallel computation
with nogil, parallel(num_threads=num_threads):
# Allocate thread-local memory
indices = ...
sorted_scores = ...
cumsum = ...
penalties_cumsum = ...
L = ...
# Pre-compute penalty terms for different set sizes
for k in prange(k_star, n_classes):
penalties_cumsum[k] = lambda_star * (k - k_star + 1)
# Main parallel loop over all samples
for i in prange(n_samples):
cdef intp_t inc = i * n_classes
# Sort prediction probabilities and store their original indices
_argsort(indices + inc, &oob_pred[i, 0], n_classes)
# Compute cumulative sum of sorted probabilities
sorted_scores[inc] = oob_pred[i, indices[inc]]
cumsum[inc] = sorted_scores[inc]
for j in range(1, n_classes):
sorted_scores[inc+j] = oob_pred[i, indices[inc+j]]
cumsum[inc+j] = cumsum[inc+j-1]
+ sorted_scores[inc+j]
# Determine the size of the prediction set (L)
L[i] = 1
for j in range(n_classes):
if (cumsum[inc+j] + penalties_cumsum[j]) <= tau:
L[i] = j + 2
L[i] = min(L[i], n_classes)
# Optional randomization step to potentially shrink the set size by one
if randomized:
for i in prange(n_samples):
# U[i] ~ Uniform([0, 1]).
# For sufficiently large U[i],
L[i] = L[i] - 1
# Construct the final prediction sets (binary matrix)
for i in prange(n_samples):
inc = i * n_classes
for j in range(L[i]):
out[i, indices[inc+j]] = 1
For cross-conformal methods like CV+ and J+ab, conformity scores for test samples must be computed using the out-of-bag (OOB) predictions from the ensemble. The core logic of _compute_test_giqs_cv() mirrors the RAPS score calculation from the split conformal case but is applied over a rank 3 array of shape (n_train, n_test, n_classes). For each training sample $i$, the corresponding OOB model is used to generate scores for all test samples. This process is parallelized across the training samples, leading to a significant speed-up.
float64_t[:,:,::1] out,
int64_t k_star,
float32_t lambda_star,
bint randomized,
bint allow_empty_sets,
intp_t num_threads,
intp_t random_state):
"""Compute generalized inverse quantiles for CV+ and
Jackknife+-after-Bootstrap methods.
# Variable declarations
cdef Py_ssize_t i, j, k
cdef Py_ssize_t n_samples = oob_pred.shape[0]
cdef Py_ssize_t n_test = oob_pred.shape[1]
cdef Py_ssize_t n_classes = oob_pred.shape[2]
cdef int32_t* I
cdef float64_t* sorted_scores, *penalty, *E
cdef float32_t* U
# Release the GIL for parallel computation
with nogil, parallel(num_threads=num_threads):
# Allocate thread-local memory
I = ...
sorted_scores = ...
# Parallel loop over training samples
for i in prange(n_samples):
# Inner loop over test samples
for j in range(n_test):
# Sort probabilities for each test sample and compute cumulative sum
cdef intp_t ij_idx = (i*n_test+j) * n_classes
_argsort(&I[ij_idx], &oob_pred[i,j,0], n_classes)
for k in range(n_classes):
sorted_scores[ij_idx+k] = oob_pred[i,j,I[ij_idx+k]]
# Calculate the cumulative sum of probabilities
out[i,j,0] = sorted_scores[ij_idx]
for k in range(1, n_classes):
out[i,j,k] = out[i,j,k-1]
+ sorted_scores[ij_idx+k]
# Apply regularization and randomization if specified
if randomized:
# Allocate thread-local memory
penalty = ...
E = ...
# Pre-compute penalties
for k in prange(k_star, n_classes):
penalty[k] = lambda_star * (k-k_star+1)
# Parallel loop to apply randomization and penalties
for i in prange(n_samples):
for j in range(n_test):
# U[j] ~ Uniform([0,1]).
ij_idx = i*samples_stride + j*n_classes
# Calculate the RAPS score for the most
# probable class. If empty sets are not
# allowed, U is set to 1 for the class
# to be included in the prediction set.
if allow_empty_sets:
E[ij_idx] = U[j] * out[i,j,0] + penalty[0]
else:
E[ij_idx] = out[i,j,0] + penalty[0]
# Iterate through the remaining classes.
for k in range(1, n_classes):
# Calculate RAPS score for the k-th class:
E[ij_idx+k] = (U[j]*sorted_scores[ij_idx+k]
+ out[i,j,k-1] + penalty[k])
# Reorder the computed scores from rank order
# back to the original class indices
for k in range(n_classes):
out[i,j,I[ij_idx+k]] = E[ij_idx+k]
else: # Non-randomized case
for i in prange(n_samples):
for j in range(n_test):
for k in range(k_star, n_classes):
out[i,j,k] += lambda_star